# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

### v1

@@ Abstract
The Poisson compound decision problem is a classical problem in statistics, for which parametric and nonparametric empirical Bayes methodologies are available to estimate the Poisson’s means in static or batch domains. In this paper, we consider the Poisson compound decision problem in a streaming or online domain. By relying on a quasi-Bayesian approach, often referred to as Newton’s algorithm, we obtain sequential Poisson’s mean estimates that are of easy evaluation, computationally efficient and with a constant computational cost as data increase, which is desirable for streaming data. Large sample asymptotic properties of the proposed estimates are investigated, also providing frequentist guarantees in terms of a regret analysis. We validate empirically our methodology, both on synthetic and real data, comparing against the most popular alternatives.
Keywords: Empirical Bayes; Newton’s algorithm; Poisson compound decision; quasi-Bayes; streaming data.
@@ 1 Introduction
Given $n\geq 1$ observations modeled as independent Poisson random variables $Y_{1},\ldots,Y_{n}$ with means $\theta_{1},\ldots,\theta_{n}$, respectively, the Poisson compound decision problem refers to the estimation of $(\theta_{1},\ldots,\theta_{n})$ under the squared-error loss. Empirical Bayes is a general approach to handle compound decision problems as first suggested by Robbins (1951, 1956); see also Efron (2003) and Zhang (2003). Assuming the parameters $\theta_{i}$’s to be i.i.d as a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$, the best estimate of $\theta_{i}$ under the squared-error loss, i.e. the Bayes estimate, is the posterior mean, namely for $y\in\mathbb{N}_{0}$
@@
where $p_{G}$ is the probability mass function of the $Y_{i}$’s. The empirical Bayes approach proposes to bypass the knowledge of $G$, which might be unavailable in practice, estimating either $p_{G}$ or ${G}$. Methodologies to estimate $p_{G}$ and ${G}$ are referred to as $f$-modeling and $g$-modeling, respectively (Efron, 2014).
Under $f$-modeling, one obtains an estimate of $p_{G}$ from $(Y_{1},\ldots,Y_{n})$, and then apply (1). For example, Robbins (1956) proposed an estimate of $\hat{\theta}_{G}(y)$ by replacing $p_{G}$ in (1) with the empirical distribution
$\hat{p}_{G}(y)=\sum_{i=1}^{n}I(Y_{i}=y)\qquad y\in\mathbb{N}_{0},$ (2)
where $I$ denotes the indicator function. This is the celebrated nonparametric empirical Bayes estimate (Brown et al., 2013; Polyanskiy and Wu, 2021). Despite its simplicity and (minimax) optimality, Robbins’ method produces unstable estimates in practice; this occurs particularly when $\hat{p}_{G}$ is small or zero (Efron and Hastie, 2021). To alleviate the shortcomings of Robbins’ method, a number of modifications that enforce smoothness of $\hat{p}_{G}(y)$ have been proposed, including $g$-modeling.
Under $g$-modeling, one obtains an estimate of $G$ from $(Y_{1},\ldots,Y_{n})$, and then apply (1). An example is the estimate of $\hat{\theta}_{G}(y)$ obtained by replacing $G$ in (1) with the nonparametric maximum likelihood estimate
$\hat{G}=\operatorname*{arg\,max}_{G}\frac{1}{n}\sum_{i=1}^{n}\log p_{G}(Y_{i}),$ (3)
where the maximization is taken over all distributions on $\mathbb{R}^{+}$ (Kiefer and Wolfowitz, 1956). Much of $g$-modeling involves the estimate (3) (Laird, 1978; Jiang and Zhang, 2009) and nonparametric minimum distance estimates (Jana et al., 2024). Parametric $g$-modeling has been much less developed (Efron and Morris, 1978; Efron, 2016). A fully Bayesian approach is possible by placing a hyperprior distribution on $G$; this is the Bayes empirical Bayes method (Deely and Lindley, 1981).
@@
The Poisson compound decision problem provides a flexible framework for count data, with a broad range of applications where static or batch processing have traditionally been sufficient, e.g. missing-species problems and microarrays (Efron, 2003), insurance analytics (Efron and Hastie, 2021, Chapter 6), deconvolution (Efron, 2016), data confidentiality (Zhang, 2005) and sports analytics (Jana et al., 2024). However, the emergence of streaming count data in some of these applications, as well as in social media analytics, network traffic monitoring, bio-surveillance and healthcare, and e-commerce web analytics, makes it natural and necessary to consider the Poisson compound decision problem in a streaming or online domain, where a timely analysis can enhance decision-making and responsiveness to new information. While it is indeed possible to adapt existing empirical Bayes methodologies to streaming data, the resulting estimates typically suffer from computational inefficiencies, especially when they involve optimizations and posterior computations.
1.1 Our contribution
@@
We propose a principled sequential $g$-modeling methodology for the Poisson compound decision problem: starting from an initial guess $G_{0}$ for the prior $G$, for $n\geq 1$ we update $G_{n-1}$ at each new observation $Y_{n}$ following a procedure of Smith and Makov (1978), often referred to as Newton’s algorithm (Newton et al., 1998; Martin and Ghosh, 2008). This recursive procedure is however regarded not just as an algorithm, but as a quasi-Bayesian learning model (Fortini and Petrone, 2020), which provides a predictive construction for an asymptotically equivalent Bayesian model. The resulting quasi-Bayes $g$-modeling allows to estimate $\hat{\theta}_{G}(y)$, based on $(Y_{1},\dots,Y_{n})$, with the plug-in estimate $\hat{\theta}_{G_{n}}(y)$ obtained from (1) by replacing $G$ with $G_{n}$. Besides its ease of evaluation, the sequential estimate $\hat{\theta}_{G_{n}}(y)$ is computationally efficient and with constant computational cost as data increase, which is desirable for streaming data. By considering Newton’s algorithm as a learning model, it is possible to measure uncertainty of $\hat{\theta}_{G_{n}}(y)$; here, we obtain a Gaussian central limit theorem for $\hat{\theta}_{G_{n}}(y)$, $n\rightarrow+\infty$, leading to asymptotic credible intervals for $\hat{\theta}_{G}(y)$.
@@
We provide the estimate $\hat{\theta}_{G_{n}}(y)$ with asymptotic frequentist guarantees. Assuming the “true” mixing distribution $G$ to be known, say equal to the oracle prior $G^{\ast}$, and that the observations $Y_{n}$’s are i.i.d. from a Poisson mixture model with mixing distribution $G^{*}$, we study the large $n$ asymptotic behavior of the Bayesian regret of $\hat{\theta}_{G_{n}}(y)$, which is defined as the difference between the Bayes risk of $\hat{\theta}_{G_{n}}(y)$ and the Bayes risk of the plug-in estimate $\hat{\theta}_{G^{\ast}}(y)$ obtained from (1) by replacing $G$ with $G^{\ast}$. In particular, we characterize the rate at which the Bayesian regret goes to zero as $n\rightarrow+\infty$.
@@
An empirical validation of our sequential $g$-modeling is presented on synthetic data generated i.i.d from a “true” mixing distribution $G$, and on real data. The estimate $\hat{\theta}_{G_{n}}(y)$, at fixed $n\geq 1$, outperforms the nonparametric empirical Bayes method and competes with the nonparametric maximum likelihood and the minimum squared Hellinger distance methods. Further, $\hat{\theta}_{G_{n}}(y)$ competes with the parametric Bayes empirical Bayes method, assuming a Gamma hyperprior on $G$.
@@ 2 Quasi-Bayes empirical Bayes
We consider a stream of observations $(Y_{n})_{n\geq 1}$, such that the $Y_{n}$’s have Poisson distributions with means $\theta_{n}$’s, respectively. To estimate sequentially the parameters $\theta_{n}$’s, we adopt a $g$-modeling methodology that, on a probability space $(\Omega,\mathcal{F},\mathds{P})$, specifies sequentially the distribution of $(Y_{n})_{n\geq 1}$ by giving the conditional distribution of $Y_{n+1}$ given $(Y_{1},\dots,Y_{n})$, for each $n\geq 0$. In particular, denoting by $(\mathcal{G}_{n})_{n\geq 1}$ the natural filtration of $(Y_{n})_{n\geq 1}$, for $n\geq 0$ each conditional distribution with respect to $\mathcal{G}_{n}$ is assumed to be a mixture of Poisson kernels $k(\cdot\mid\theta)$, with $\theta$ defined in a sufficiently fine and wide grid of points $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, with $\vartheta_{j}<\vartheta_{j+1}$ ($j=1,\dots,d-1$), namely
@@
given an initial guess $g_{0}$ and a sequence $(\alpha_{n})_{n\geq 1}$ of real numbers in $(0,1)$ such that $\sum_{n\geq 1}\alpha_{n}=+\infty$ and $\sum_{n\geq 1}\alpha^{2}_{n}<+\infty$. From (5), $g_{n}$ is updated as $Y_{n+1}$ becomes available, by taking a weighted average, with respect to $\alpha_{n}$, of $g_{n}$ itself and of its posterior distribution based on $Y_{n+1}$. The sequence $(\alpha_{n})_{n\geq 1}$ is referred to as the learning rate, for which a standard choice is $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$. The next result explains in which sense our sequential $g$-modeling is quasi-Bayesian.
@@
See Supplementary Material C.1 for the proof of Theorem 2.1. Under (4)-(5), for each $n\geq 0$, the future observations of $(Y_{n})_{n\geq 1}$ are conditionally identically distributed given $\mathcal{G}_{n}$. From Berti et al. (2004), this implies that the $Y_{n}$’s are asymptotically i.i.d., given $\tilde{g}$, with law $p_{\tilde{g}}$, taking on the interpretation of a quasi-Bayesian framework (Fortini and Petrone, 2020). In particular, $p_{\tilde{g}}$ is the random distribution that would be assigned if we had access to an infinitely large sample, from which it is reasonable to consider $\tilde{g}$ as the parameter driving our sequential $g$-modeling. Differently from the quasi-Bayesian framework, the Bayesian framework relies on a prior distribution on $g$, such that the $Y_{n}$’s are i.i.d, given $g$. Under this framework, the computation cost to evaluate the posterior distribution of $g$ is typically cumbersome, and the posterior distribution must be recomputed for each new observation. The quasi-Bayesian framework offers a significantly lower computational cost.
2.1 Sequential estimation
Under the sequential $g$-modeling methodology of Section 2, the empirical Bayes estimate of $\hat{\theta}_{G}(y)$ in (1) that would be assigned if we had access to an infinitely large sample would be of the form
@@ 2.1 Sequential estimation
An estimate of $\hat{\theta}_{G}(y)$ based on the sample $(Y_{1},\ldots,Y_{n})$ is obtained from (6) by replacing $\tilde{g}$ with $g_{n}$, i.e.
$\hat{\theta}_{g_{n}}(y)=(y+1)\frac{p_{g_{n}}(y+1)}{p_{g_{n}}(y)}\qquad y\in\mathbb{N}_{0}.$ (7)
Although the mixing distribution is concentrated on $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, we are not restricting the parameter $\theta$ or its estimate to take values in $\Theta_{d}$. Under (4)-(5), $g_{n}$ converges $\mathds{P}$ a.s. to $\tilde{g}$, as $n\rightarrow+\infty$. Then, it follows that for every $y\in\mathbb{N}_{0}$, $\hat{\theta}_{g_{n}}(y)$ converges $\mathds{P}$ a.s. to $\hat{\theta}_{\tilde{g}}(y)$, as $n\rightarrow+\infty$. Next theorem gives an asymptotic Gaussian approximation of the posterior distribution of $\hat{\theta}_{\tilde{g}}(y)$ for $y\in\mathbb{N}_{0}$.
@@
See Supplementary Material C.2 for the proof of Theorem 2.2. For the standard choice $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$ then $b_{n}=(2\gamma-1)n^{2\gamma-1}$. Let $W_{g_{n}}(y)$ be defined as in (8), with $g_{n}$ in the place of $\tilde{g}$. Since $g_{n}$ converges to $\tilde{g}$ ($\mathds{P}$-a.s.), then $W_{g_{n}}(y)$ converges to $W_{\tilde{g}}(y)$, $\mathds{P}$-a.s. Therefore, the conditional distribution of $\hat{\theta}_{\tilde{g}}(y)$ given $\mathcal{G}_{n}$ is approximately Gaussian with mean $\hat{\theta}_{g_{n}}(y)$ and variance $W_{g_{n}}(y)$. This provides asymptotic credible intervals for $\hat{\theta}_{\tilde{g}}(y)$, for $y\in\mathbb{N}_{0}$.
Remark 2.3.
A sequential $g$-modeling methodology may be also considered from a frequentist perspective, by assuming i.i.d. observations $(Y_{1},\ldots,Y_{n})$ and applying Newton’s algorithm as a consistent estimate of a “true” $g$ (Tokdar et al., 2009; Martin and Tokdar, 2009). This approach would provide the same estimate $\hat{\theta}_{g_{n}}(y)$ of $\hat{\theta}_{G}(y)$ as in (7), and an upper bound for the rate of convergence (see Section 2.2). However, to the best of our knowledge, going beyond point estimation, for example by building asymptotic confidence intervals, is not feasible based on asymptotic results available for Newton’s algorithm. Indeed, these are based on stochastic approximation (Martin and Ghosh, 2008), but central limit theorems in this field does not apply to Newton’s algorithm.

### v2

@@ Abstract
The Poisson compound decision problem is a long-standing problem in statistics, where empirical Bayes methodologies are commonly used to estimate Poisson’s means in static or batch domains. In this paper, we study the Poisson compound decision problem in a streaming or online domain. Adopting a quasi-Bayesian approach, referred to as Newton’s algorithm, we obtain a sequential estimate that is easy to evaluate, computationally efficient, and maintain a constant per-observation computational cost as data accumulate. Asymptotic frequentist guarantees of this estimate are established, showing consistency and asymptotic optimality, where the latter is understood as vanishing excess Bayes risk or regret. We demonstrate the effectiveness of our methodology through empirical analysis on synthetic and real data, with comparisons to existing approaches.
Keywords: Empirical Bayes; Newton’s algorithm; Poisson compound decision; quasi-Bayes; streaming data
@@ 1 Introduction
Given $n\geq 1$ observations modeled as independent Poisson random variables $Y_{1},\ldots,Y_{n}$ with means $\theta_{1},\ldots,\theta_{n}$, respectively, the Poisson compound decision problem refers to the estimation of $(\theta_{1},\ldots,\theta_{n})$ under the squared-error loss. Empirical Bayes is a general approach to handle compound decision problems (Robbins, 1951, 1956; Zhang, 2003). Assuming the $\theta_{i}$’s to be i.i.d. as a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$, the best estimate of $\theta_{i}$ is the posterior mean or Bayes estimate, namely for $y\in\mathbb{N}_{0}$
@@
where $p_{G}$ is the probability mass function of the $Y_{i}$’s. The empirical Bayes approach proposes to bypass the knowledge of $G$, which might be unavailable in practice, estimating either the marginal $p_{G}$, a strategy known as $f$-modeling, or the prior ${G}$, a strategy known as $g$-modeling (Efron, 2014).
1.1 Background and motivation
A prominent example of $f$-modeling is Robbins’ method (Robbins, 1956), which estimates $p_{G}$ in (1) with the empirical distribution of the $Y_{i}$’s (Brown et al., 2013; Polyanskiy and Wu, 2021; Jana et al., 2023); see Efron (2014, Section 3 and Section 4). Although widely used in practice for its computational simplicity, $f$-modeling lacks robustness and numerical stability; in fact, it is easily derailed by “outliers” in the data, namely data points that appear only a few times, causing estimators to take exceptionally small or large values (Efron and Hastie, 2021; Shen and Wu, 2024).
Examples of $g$-modeling include estimates of $G$ in (1) based on maximum likelihood and minimum distance methods (Jiang and Zhang, 2009; Jana et al., 2024); see Efron (2014, Section 5). A fully Bayesian approach is also possible by placing a hyperprior on $G$, a strategy known as Bayes empirical Bayes (Deely and Lindley, 1981; Efron, 2019). While $g$-modeling yields more accurate estimates than $f$-modeling, as well as better incorporating prior knowledge, it is computationally demanding, particularly in nonparametric and high-dimensional settings (Jana et al., 2023).
@@ 1.1 Background and motivation
The Poisson compound decision problem offers a flexible framework for modeling count data, with successful applications in settings where static or batch processing has traditionally sufficed, such as microarrays (Efron, 2003), insurance analytics (Efron and Hastie, 2021, Chapter 6), deconvolution (Efron, 2016), data confidentiality (Zhang, 2005) and sports analytics (Jana et al., 2024). However, the rise of streaming data in some of these fields, as well as in social media analytics, network traffic monitoring, bio-surveillance and healthcare, and e-commerce web analytics, calls for extending the Poisson compound decision problem to the online setting, where timely analysis is critical for responsive and informed decision-making. Although the existing empirical Bayes methods can, in principle, be adapted to streaming data, they often suffer from computational inefficiencies, particularly when involving optimization or posterior computations.
1.2 Preview of our contributions
@@
We propose a principled sequential $g$-modeling methodology for the Poisson compound decision problem. Starting from an initial guess $G_{0}$ for $G$, for $n\geq 1$ our approach updates $G_{n-1}$ at each new observation $Y_{n}$ following a procedure of Smith and Makov (1978), referred to as Newton’s algorithm (Newton et al., 1998; Martin and Ghosh, 2008). This recursive procedure is regarded as a quasi-Bayesian learning model, yielding a predictive construction for a large-$n$ asymptotically equivalent Bayesian model (Fortini and Petrone, 2020). A quasi-Bayes empirical Bayes estimate of $\hat{\theta}_{G}(y)$ is then obtained from (1) by replacing $G$ with $G_{n}$. Besides being easy to evaluate, the estimate $\hat{\theta}_{G_{n}}(y)$ is computationally efficient, and maintain a constant per-observation computational cost as data accumulate. Considering Newton’s algorithm as a learning model allows to measure uncertainty for $\hat{\theta}_{G_{n}}(y)$, for which we build an asymptotic credible interval via the central limit theorem.
@@
Asymptotic frequentist guarantees are established for $\hat{\theta}_{G_{n}}(y)$, assuming that the observations $Y_{n}$’s are i.i.d. from a Poisson mixture model with a “true” mixing distribution, say the oracle prior $G^{\ast}$. We provide sufficient conditions on $G^{\ast}$ to ensure consistency and asymptotic optimality of $\hat{\theta}_{G_{n}}(y)$ as $n\rightarrow+\infty$ , where the latter is understood as vanishing excess Bayes risk or regret. Under additional assumptions on Newton’s algorithm, the rate at which the regret converges to zero is also derived.
@@
We validate our methodology empirically on synthetic and real data. For fixed $n\geq 1$, the estimate $\hat{\theta}_{G_{n}}(y)$ consistently outperforms Robbins’ method and performs competitively with the nonparametric maximum likelihood method and the minimum squared Hellinger distance method. It also performs on par with a parametric Bayes empirical Bayes method with a Gamma hyperprior.
@@ 2 Quasi-Bayes empirical Bayes
For a stream of observations $(Y_{n})_{n\geq 1}$, we propose a sequential $g$-modeling methodology that, on a probability space $(\Omega,\mathcal{F},\mathds{P})$, specifies the distribution of $(Y_{n})_{n\geq 1}$ through the conditional distribution of $Y_{n+1}$ given $(Y_{1},\dots,Y_{n})$, for each $n\geq 0$. As in the Bayesian framework, the model is conceived here as a learning model and not as a data-generating process (Fortini and Petrone, 2025).
Letting $(\mathcal{G}_{n})_{n\geq 1}$ denote the natural filtration of $(Y_{n})_{n\geq 1}$, for each $n\geq 0$ the conditional distribution of $Y_{n+1}$ given $\mathcal{G}_{n}$ is modeled according to a mixture of the Poisson kernel $k(\cdot\mid\theta)$, with $\theta$ ranging over a sufficiently fine and wide grid $\Theta_{d}=\{\vartheta_{1},\dots,\vartheta_{d}\}$, with $\vartheta_{j}<\vartheta_{j+1}$ for $j=1,\dots,d-1$, i.e.,
@@
given an initial guess $g_{0}$ and the $\alpha_{n}$’s in $(0,1)$ such that $\sum_{n\geq 1}\alpha_{n}=+\infty$ and $\sum_{n\geq 1}\alpha^{2}_{n}<+\infty$. From (3), the learning model updates $g_{n}$ as $Y_{n+1}$ becomes available, by taking a weighted average, with respect to $\alpha_{n}$, between $g_{n}$ and its posterior distribution based on $Y_{n+1}$. The sequence $(\alpha_{n})_{n\geq 1}$ is referred to as the learning rate, for which a standard choice is $\alpha_{n}=(\alpha+n)^{-\gamma}$ for $\alpha>0$ and $\gamma\in(1/2,1]$.
The following proposition outlines an effective procedure for constructing the grid $\Theta_{d}$ used in (2)-(3). Beyond serving as a practical tool for implementing the sequential $g$-modeling methodology, this construction also plays a central role in establishing asymptotic frequentist guarantees. We denote by $\text{KL}(p||q)$ the Kullback-Leibler divergence between the probability distributions $p$ and $q$.
Proposition 2.1.
Let the prior $G$ in (1) be such that $\sum_{y\in\mathbb{N}_{0}}y^{k}p_{G}(y)\leq m_{k}$ for some $k\in\mathbb{N}$ with $k\geq 2$ and $m_{k}\in\mathbb{R}$, and for a fixed $\eta>0$ define the equally spaced grid $\Theta_{\eta}=\{\vartheta_{i}=i\eta\text{ : }i=1,\ldots,d_{\eta}\}$ with $d_{\eta}=\inf\{n\in\mathbb{N}\text{ : }n>\eta^{-1}\text{ and }n^{1-k}\log(n\eta)m_{k}\leq\eta^{k}\}$. If $\mathcal{S}_{\eta}$ is the set of all the probability mass functions on $\Theta_{\eta}$, then $g_{\eta}=\text{argmin}_{f\in\mathcal{S}_{\eta}}\text{KL}(p_{G}||p_{f})$ is unique and $\text{KL}(p_{G}||p_{g_{\eta}})<2\eta$.
See C.1 for the proof of Proposition 2.1. The probability mass function $g_{\eta}$ in Proposition 2.1 is referred to as the $d_{\eta}$-KL-discretization of a prior $G$ on $\Theta\subseteq\mathbb{R}^{+}$. This form of discretization is of independent interest in $g$-modeling approaches to the Poisson compound decision problem (Jana et al., 2024).
2.1 Quasi-Bayes properties
We show that learning model defined by (2)-(3) is quasi-Bayesian in the sense of Fortini and Petrone (2020), yielding a predictive construction for a large-$n$ asymptotically equivalent Bayesian model.
@@ 2.1 Quasi-Bayes properties
See C.2 for the proof of Theorem 2.2. Under the learning model (2)-(3), for each $n\geq 0$, the future observations of $(Y_{n})_{n\geq 1}$ are considered conditionally identically distributed given $\mathcal{G}_{n}$. From Berti et al. (2004, Theorem 2.2), this implies that the $Y_{n}$’s are asymptotically i.i.d., given the random limit $\tilde{g}$, with law $p_{\tilde{g}}$, which fits within the quasi-Bayesian framework of Fortini and Petrone (2020). In this perspective, $p_{\tilde{g}}$ represents the random distribution that would be assigned if we had access to an infinitely large sample, which makes $\tilde{g}$ the parameter of interest. Accordingly, (1) becomes
@@
Differently from the Bayesian framework, which posits a prior distribution on $g$ such that the $Y_{n}$’s are i.i.d. given $g$, the quasi-Bayesian approach avoids the computational cost of recomputing the posterior distribution with each new observation, thereby offering substantial computational efficiency.
2.2 Sequential estimation and credible intervals
Under the learning model (2)-(3), an estimate of $\hat{\theta}_{\tilde{g}}(y)$ is obtained by replacing $\tilde{g}$ in (4) with $g_{n}$, i.e.
$\hat{\theta}_{g_{n}}(y)=(y+1)\frac{p_{g_{n}}(y+1)}{p_{g_{n}}(y)}\qquad y\in\mathbb{N}_{0},$ (5)
which is referred to as quasi-Bayes empirical Bayes. As a corollary of Theorem 2.2, for every $y\in\mathbb{N}_{0}$ it holds that $\hat{\theta}_{g_{n}}(y)$ converges $\mathds{P}$ a.s. to $\hat{\theta}_{\tilde{g}}(y)$, as $n\rightarrow+\infty$. The following central limit theorem allows to measure uncertainty for $\hat{\theta}_{G_{n}}(y)$, building an asymptotic credible interval for $\hat{\theta}_{\tilde{g}}(y)$.
@@ 2.2 Sequential estimation and credible intervals
See C.3 for the proof of Theorem 2.3. If $W_{g_{n}}(y)$ is defined as in (6), with $g_{n}$ in the place of $\tilde{g}$, then for every $y\in\mathbb{N}_{0}$ it holds that $W_{g_{n}}(y)$ converges $\mathds{P}$-a.s. to $W_{\tilde{g}}(y)$, as $n\rightarrow+\infty$. Therefore, the conditional distribution of $\hat{\theta}_{\tilde{g}}(y)$ given $\mathcal{G}_{n}$ is approximately a Gaussian distribution with mean $\hat{\theta}_{g_{n}}(y)$ and variance $W_{g_{n}}(y)$. This result provides an asymptotic credible interval for $\hat{\theta}_{\tilde{g}}(y)$, for $y\in\mathbb{N}_{0}$.
A sequential $g$-modeling methodology may be also considered from a frequentist perspective, where $(Y_{n})_{n\geq 1}$ are i.i.d. from a Poisson mixture model with a “true” mixing distribution $g^{\ast}$, and Newton’s algorithm is employed as a consistent estimate of $g^{\ast}$ (Tokdar et al., 2009; Martin and Tokdar, 2009). While this perspective leads to the same estimate $\hat{\theta}_{g_{n}}(y)$ as (5), constructing asymptotic confidence intervals remains out of reach, as the stochastic approximation framework underlying Newton’s method (Martin and Ghosh, 2008) does not apply to Gaussian central limit theorems.